# Bridging Developer Needs and Feasible Features for AI Assistants in IDEs

### v2

@@ IV-A1 Materials
Based on insights from interviews, we generated an initial list of 38 potential AI-based IDE features [28]. This list was reviewed by seven experienced researchers from industry and academia: three specialists in human–AI experience, one assistant professor in computer science, and three senior software engineering researchers. All reviewers had more than five years of professional experience. Through this collaborative process, we selected 12 features for the prediction market. These were chosen based on their perceived market feasibility within the next year and the possibility of defining each feature with a high degree of clarity.
@@
Based on the final list of features, we formulated 24 binary questions for the prediction market. Each question asked whether a given AI-based feature would be released within the next six months, either in JetBrains IDEs or any other AI-enabled development tool. These questions were published on JetBrains ’ internal prediction market platform, where employees place virtual bets to estimate the likelihood of future events. All 24 questions were grouped into a single tournament, where cash prizes (\$1500, \$1000, \$500) were to be awarded to the top three predictors on the leader board to incentivize participation.
IV-A2 Participants
@@
All participants of the prediction market were employed by JetBrains. A total of 102 participants [28] across 84 teams placed bets on at least one question during the tournament. The median tenure was 4 years and 10 months. The most common roles included Software Developer (34 participants), Team Lead (20), Product Manager (6), Analyst (6), QA Engineer (5), and ML Engineer (4).
IV-A3 Data collection
@@
The tournament was launched in December 2024. Each participant received 500 virtual coins to place bets on the outcomes of 24 binary questions. For each question, participants could allocate any number of coins to the YES or NO outcome. Before trading, all participants were informed about tournament rules, trading procedures, and strategies for making informed bets. During trading, the current market probability for each question was visible. Bidding closed on June 1, 2025, and all questions were resolved in July 2025. Participants who made correct predictions earned coins, and the top performers were awarded monetary prizes.
@@
The final resolutions were made by a panel of four expert judges, each with over seven years of professional experience in software engineering, AI/ML research, and product development at the company. The panel of judges did not participate in the initial preparation of the features for prediction. Before resolution, the organizing team compiled a summary report [28] on the market status of each feature based on analysis of the feature set of JetBrains’ IDEs, and 20 other tools that were considered, e.g., Windsurf, Cursor, GitHub Copilot. This document served as a reference for the judges, who reached each final decision through several rounds of structured discussion.
IV-A4 Data analysis
@@
We analyzed the prediction market results along three dimensions: participant engagement, accuracy of forecasts, and final probability estimates. Engagement was measured using two metrics: trading volume and market liquidity. Trading Volume captures the total number of virtual coins bet across both outcomes of a question, reflecting overall activity. Market Liquidity reflects the breadth of participation and was calculated by assigning a base value of 50 to each question, with 20 additional points added for every unique participant who placed a bet. Accuracy was assessed by comparing the community’s majority prediction with the actual outcome, as defined by the resolution criteria. The final probability estimate for each question was calculated as the proportion of coins placed on the YES outcome relative to the total of coins bet on that question.
IV-B FINDINGS
@@
To assess whether internal practitioners recognize and correctly anticipate which AI features inspired by user needs are likely to be implemented, we analyzed prediction market outcomes along three dimensions: forecast accuracy, final probability estimates, and participant engagement. The results are summarized in Table I.
@@
Ultimately, experts concluded that 6 of the 12 features have appeared on the market in any coding tool, and 2 of the 12 appeared specifically in JetBrains IDEs. The market produced accurate forecasts for 20 of the 24 questions (83%). Accuracy was higher for internal JetBrains releases (92%) than for external tooling (75%), suggesting better visibility into internal development plans and greater uncertainty around the broader field.
All four incorrect forecasts were false negatives: participants assigned low probabilities to features that were actually released. These included Project-wide Context Awareness in JetBrains IDEs, and three features that appeared in other tools: Annotating AI-Generated Code, Business-context Awareness, and the Daily Ecosystem Summarizer (provides actionable summaries of workspace activities to support informed decisions). This pattern reveals a blind spot: while these needs were voiced by users, they were underestimated by internal practitioners, potentially due to organizational constraints or differing views on feasibility. What is interesting, no feature was overestimated, meaning there were no cases where the market predicted that a feature would be implemented, but it actually was not. This conservative bias suggests that market signals can be trusted to filter out overly speculative ideas (e.g., Brain-computer Interface Coding), but may understate the pace of external innovation.
@@
Among the 12 AI features derived from user needs in Phase 1, three were correctly forecast as likely to be implemented and were in fact released: Multi-file Generation, Project-wide Context Awareness, and Sketch2Code. These features respond to the needs around AI in Implementation and Context-aware Technology. Practitioners identified them as feasible and relevant, and their accurate prediction indicates strong consensus and implementation readiness.
In contrast, features that required deeper system integration or organizational data, such as Proactive Documentation, Automation of Repetitive Tasks, Debugging, Optimization, and AI Real-Time Refactoring, received low probability scores (all below 33%) and were not released. This suggests that despite user interest, practitioners recognized the technical or infrastructural barriers involved. At the same time, the low probabilities assigned to these features do not imply irrelevance. On the contrary, their presence in the prioritized list from Phase 1 and the fact that practitioners recognized them, albeit cautiously, indicates that these needs remain unaddressed by current tools. This creates a clear opportunity for future development.
@@
Answer to RQ3. Practitioners correctly predicted the outcome for 20 of the 24 features, showing that collective judgment aligns with both user needs and actual realization. Features related to implementation and context awareness were accurately anticipated and released. In contrast, features involving proactivity and code maintenance were consistently underestimated, indicating that these needs remain unmet.
V DISCUSSION
@@
While past studies have mapped general expectations for AI tools for coding, our work focuses on in-IDE use, capturing and systematizing nuanced needs shaped by the real-time and context-rich nature of coding environments. Our findings support existing calls for context-aware [14], non-intrusive [32, 17], transparent [33], ethical [13, 12], and customizable [11] AI systems. However, our study also reveals new priorities: integration into broader business workflows, support for skill building, and proactive behavior.
The inclusion of Churners and Non-Users helps surface barriers often missed in adoption-centered research [7]. In particular, there is a need for trust-building mechanisms and accessible onboarding.
@@
Create Proactive AI. All three participant groups expressed a need for Proactive AI, which was also rated as relevant in the prediction market. AI assistants should take a more proactive role in development, moving beyond reactive responses to explicit prompts and recommendations. Key actions include:
  • Autonomous suggestions. Continuously analyze the code and suggest improvements, such as identifying potential bugs, performance issues, or optimization opportunities.
  • Automated actions. Instead of suggesting changes, the assistant should be able to perform them proactively, e.g., automatically creating new files with boilerplate code or project structure when a user starts a new project.
Build User Trust. Trust is critical for AI adoption in development environments. Developers should prioritize features that build transparency and reliability:
  • On-premise AI. Offer on-premise AI to enhance security and privacy for organizations with sensitive codebases.
  • Transparent outputs. Providing explanations and sources (e.g., Stack Overflow links) for the assistant’s suggestions helps users understand and trust the output.
  • Up-to-date data. Ensure that the AI assistant leverages current and relevant libraries, frameworks, and coding standards to provide reliable recommendations.
Ensure Accessibility. AI technologies should be usable by developers from various regions and backgrounds:
  • Flexible pricing and access. Enable users to choose the features they want to pay for and provide options to access AI from various platforms.
  • Language support. Incorporate multi-language support and localized interfaces to ensure that AI tools are usable in different countries.
  • Inclusive access. Consider regional differences in infrastructure by offering lightweight or offline versions for areas with limited internet connectivity.
Ensure Privacy. Privacy and security were unresolved concerns for many participants. To support adoption under organizational and national constraints, AI systems should:
@@
  • Support on-premise and local AI. Allow use without sending code to external servers, enabling companies to keep data within national or internal infrastructure.
  • Enable auditability. Ensure that AI-generated code is clearly marked to support inspection and maintain trust.
Simplifying Programming Tasks
@@
Look Beyond Code Writing. To support the needs of all users’ groups, AI should not only assist in code writing but also in code-base navigation, evaluation, and in improving code quality:
  • Codebase navigation. Support developers in understanding and exploring the codebase through contextual summaries and links between related components.
  • Project analysis. AI assistants should be able to provide evaluations of the codebase and suggest optimizations for performance, security, or maintainability.
Integrate AI Systems. Participants in all three groups expressed an unmet need for in-IDE AI systems that are tightly connected with internal systems and other engineering tools. Smooth integration with external systems will boost the AI’s utility throughout the SDLC:
  • CI/CD integration. Enable the AI to support continuous integration and continuous deployment pipelines, automating build and deployment tasks.
  • Task tracking integration. Connect with task management tools, allowing the AI to monitor issue statuses, update progress, and link IDE activities with external sources.
  • Commit management. Automate and manage commits, generating commit messages and simplifying the version control process within IDE.
  • Code review support. Integrate with code review platforms to assist in reviewing code, and suggesting improvements without leaving the IDE.
Technology Interaction
Streamline In-IDE Interaction. Participants emphasized the importance of minimizing context switching and ensuring smooth interaction between AI and existing IDE features. To support this:
  • Enable embedded interfaces. Keep AI interaction within the IDE through integrated chat and in-line code explanations to reduce disruption.
  • Support multimodal input. Allow voice commands or device-based control (e.g., wearables) to increase accessibility and flexibility.
Technology Customization
Personalize AI. Developers, especially Churners, seek personalized AI experiences that adapt to their unique needs and workflows. This can be achieved by:
  • Custom setups. Allow users to configure the AI assistant, e.g., selecting models, defining interaction methods, or choosing the technology stack.
  • Learning from interactions. Let the assistant learn from user feedback (e.g., declining specific completions) and adjust its behavior over time.
Simplifying Skill Building
@@
Reduce the Learning Curve. Onboarding and education are key for users to fully utilize AI assistants. All three groups of participants expressed challenges and needs in this regard:
  • User training. Offer tutorials and examples to help users interact effectively with the AI.
  • Prompting assistance. Implement prompt suggestions and provide users with feedback on their prompts to improve their AI interaction skills.
VI THREATS TO VALIDITY
@@
External validity. The sample was recruited from a pool of developers using JetBrains products. By focusing on JetBrains’ users but allowing for different AI tools, we aimed to balance consistency with a broad range of experiences, ensuring that our findings are applicable beyond any single tool. However, it may not fully capture the diversity of experiences and perspectives of developers using different IDEs, potentially limiting the generalizability of the findings.
Construct validity. Thematic analysis involved iterative coding and discussion sessions among the authors to refine categories and resolve disagreements. Multiple analysts reviewed and coded the data, which helps to ensure consistency and reliability, but some variability in interpretation is possible.
@@
In the second phase, conceptual features derived from interviews were used to formulate prediction market questions. However, these features did not always align well with concrete implementation outcomes. Some definitions remained necessarily broad, which introduced ambiguity into the resolution process. While a structured resolution guide and expert panel were used to mitigate this issue, some interpretive variability may remain, affecting the alignment between predicted outcomes and actual market developments.
VII CONCLUSION
@@
This study examined what developers want and expect from AI in IDEs, focusing on unmet needs and anticipated features. Through interviews with 35 Adopters, Churners, and Non-Users, we surfaced a structured landscape of user expectations across five areas: Technology Improvements, Programming Tasks, Interaction, Customization, and Skill Building. These highlight the demand for more reliable, integrated, and personalized AI experiences within development environments. To assess implementation feasibility, we conducted an internal prediction market with 102 practitioners betting on 12 AI features covering unmet needs. The results show strong alignment between user needs and practitioner judgment for core implementation and context-aware features. However, proactive behaviors and long-term code maintenance remain both underdelivered and underestimated, pointing to a systemic blind spot. Our findings bridge user expectations and product planning by showing where practitioner foresight aligns, or fails to align, with real-world developer needs. Future work should explore how to operationalize unmet needs that are recognized but deprioritized, and how to refine internal evaluation strategies for AI features beyond automation and into workflow-aware design.

### v3

@@ 4.1.1. Materials
Based on insights from interviews, we distilled an initial list of 38 potential AI-based IDE features (JetBrains Research, 2025) grounded in ideas and suggestions expressed by participants. This list was reviewed by seven experienced researchers from industry and academia: three specialists in human–AI experience, one assistant professor in computer science, and three senior software engineering researchers. All reviewers had more than five years of professional experience. Reviewers evaluated and discussed asynchronously each feature on three dimensions: existence in current tools, clarity of description, and suitability for prediction market assessment. Through this collaborative process, we selected 9 features for the prediction market.
These nine features were selected because they are representative of thematic groups that were mentioned by all three user groups (Adopters, Churners, and Non-Users) as both successes and needs. This cross-group presence indicates broad relevance and high potential for user benefit, which in turn justified their perceived feasibility within the short-term horizon of 6 months. In contrast, features from thematic groups such as AI Support in SDLC and Learning Support, while also appearing in both successes and needs categories, were expressed in ways that were too generic to be translated into concrete and clearly defined feature candidates. These were therefore excluded from the prediction market to ensure that only features with a precise and operational definition were included.
@@
Based on the final list of features, we formulated 18 binary questions for the prediction market. Each question asked whether a given AI-based feature would be released within the next six months, either in JetBrains’ IDEs or any other AI-enabled development tool. These questions were published on JetBrains’ internal prediction market platform, where employees place virtual bets to estimate the likelihood of future events. All 18 questions were grouped into a single tournament, where cash prizes (\$1500, \$1000, \$500) were to be awarded to the top three predictors on the leader board to incentivize participation.
4.1.2. Participants
@@
All participants of the prediction market were employed by JetBrains. A total of 102 participants (JetBrains Research, 2025) across 84 teams placed bets on at least one question during the tournament. The median tenure was 4 years and 10 months. The most common roles included Software Developer (34 participants), Team Lead (20), Product Manager (6), Analyst (6), Quality Assurance (QA) Engineer (5), and Machine Learning (ML) Engineer (4).
4.1.3. Data collection
@@
The tournament was launched in December 2024. Each participant received 500 virtual coins to place bets on the outcomes of 18 binary questions. For each question, participants could allocate any number of coins to the YES or NO outcome. Before trading, all participants were informed about tournament rules, trading procedures, and strategies for making informed bets. During trading, the current market probability for each question was visible. Bidding closed on June 1, 2025, and all questions were resolved in July 2025. Participants who made correct predictions earned coins, and the top performers were awarded monetary prizes.
@@
The final resolutions were made by a panel of four expert judges, each with over seven years of professional experience in software engineering, AI/ML research, and product development at the company. Judges did not participate in the initial preparation of the features for prediction. Before resolution, the organizing team compiled a summary report (JetBrains Research, 2025) on the market status of each feature based on analysis of the feature set of JetBrains’ IDEs, and 20 other tools that were considered, e.g., Windsurf, Cursor, GitHub Copilot. This document served as a reference for the judges, who reached each final decision through several rounds of structured discussion.
4.1.4. Data analysis
@@
We analyzed the prediction market results along three dimensions: participant engagement, accuracy of forecasts, and final probability estimates. Engagement was measured using two metrics: trading volume and market liquidity. Trading Volume captures the total number of virtual coins bet across both outcomes of a question, reflecting overall activity. Market Liquidity reflects the breadth of participation and was calculated by assigning a base value of 50 to each question, with 20 additional points added for every unique participant who placed a bet. Accuracy was assessed by comparing the community’s majority prediction with the actual outcome, as defined by the resolution criteria. The final probability estimate for each question was calculated as the proportion of coins placed on the YES outcome relative to the total number of coins bet on that question.
4.2. FINDINGS
@@
To assess whether internal practitioners correctly anticipate which AI features inspired by user needs are likely to be implemented, we analyzed prediction market outcomes along three dimensions: forecast accuracy, final probability estimates, and participant engagement. The results are summarized in Table 2.
@@
Ultimately, experts concluded that 4 of the 9 features have appeared on the market in any coding tool, and 2 of the 9 appeared specifically in JetBrains IDEs by the end of the prediction window (July 2025). The market produced accurate forecasts for 16 of the 18 questions (88%), showing that at least for short-term predictions the professional community demonstrates considerable precision.
Two of the incorrect forecasts were false negatives: participants assigned low probabilities to features that were actually released: Project-wide Context Awareness in JetBrains IDEs, and Business-context Awareness in any of the tools available on the market. This pattern suggests a potential a blind spot: while these needs were voiced by users, they were underestimated by internal practitioners, potentially due to organizational constraints or differing views on feasibility. What is interesting, no feature was overestimated, meaning there were no cases where the market predicted that a feature would be implemented, but it actually was not. This conservative bias suggests that market signals can be trusted to filter out overly speculative ideas (e.g., Brain-computer Interface Coding), but may understate the pace of external innovation.
@@
Among the 9 AI features derived from user needs in Phase 1, three were correctly forecast as likely to be implemented and were in fact released: Multi-file Generation, Project-wide Context Awareness, and Sketch2Code. These features respond to the needs around AI in Implementation and Context-aware Technology. Practitioners identified them as feasible and valuable, and their accurate prediction indicates strong consensus and implementation readiness.
In contrast, features that required deeper system integration or organizational data, such as Proactive Documentation, Debugging, Optimization, and AI Real-Time Refactoring, received low probability scores (all below 33%) and were not released. This suggests that despite user interest, the technical or infrastructural barriers prevent realization of such features. Here, the low probabilities assigned to these features do not imply irrelevance. On the contrary, their presence in the prioritized by experts list of interesting features from Phase 1 and the fact that practitioners did vote for them, albeit cautiously, indicates that these needs remain relevant but unaddressed by current tools. This creates a clear opportunity for future development.
@@
Future work should investigate intervention strategies to close the recognition-implementation gap. Potential approaches include: (1) capability reassessment workshops where development teams re-evaluate the feasibility of previously dismissed features using structured technical analysis; (2) incremental implementation pathways that break complex features into manageable development milestones; and (3) user-developer collaboration sessions that allow practitioners to better understand the business value and user impact of challenging-to-implement features. Additionally, longitudinal studies could track whether features initially deemed infeasible by practitioners are successfully implemented by competitor tools or through alternative technical approaches, providing evidence for or against the accuracy of initial feasibility judgments.
[…]
5. DISCUSSION
@@
While past studies have mapped general expectations for AI tools for coding, our work focuses on in-IDE use, capturing and systematizing nuanced needs shaped by the real-time and context-rich nature of coding environments. Our findings support existing calls for context-aware (Sergeyuk et al., 2024), non-intrusive (Mozannar et al., 2024; Wang et al., 2023), transparent (Zhou et al., 2024), ethical (Liu et al., 2024; Mcnutt et al., 2023), and customizable (Liang et al., 2024) AI systems. However, our study also reveals new priorities: integration into broader business workflows, support for skill building, and proactive behavior.
The inclusion of Churners and Non-Users helps surface barriers often missed in adoption-centered research (Sergeyuk et al., 2025). In particular, there is a need for trust-building mechanisms and accessible onboarding.
@@
Technology Improvements. Create Proactive AI. All three participant groups expressed a need for Proactive AI, which was also rated as relevant in the prediction market. AI assistants should take a more proactive role in development, moving beyond reactive responses to explicit prompts and recommendations. Key actions include: (1) Autonomous suggestions: continuously analyze the code and suggest improvements, such as identifying potential bugs, performance issues, or optimization opportunities; (2) Automated actions: instead of suggesting changes, the assistant should be able to perform them proactively, e.g., automatically creating new files with boilerplate code or project structure when a user starts a new project.
Build User Trust. Trust is critical for AI adoption in development environments. Developers should prioritize features that build transparency and reliability: (1) On-premise AI: offer on-premise AI to enhance security and privacy for organizations with sensitive codebases; (2) Transparent outputs: providing explanations and sources (e.g., Stack Overflow links) for the assistant’s suggestions helps users understand and trust the output; (3) Up-to-date data: ensure that the AI assistant leverages current and relevant libraries, frameworks, and coding standards to provide reliable recommendations.
Ensure Accessibility. AI technologies should be usable by developers from various regions and backgrounds: (1) Flexible pricing and access: enable users to choose the features they want to pay for and provide options to access AI from various platforms; (2) Language support: incorporate multi-language support and localized interfaces to ensure that AI tools are usable in different countries; (3) Inclusive access: consider regional differences in infrastructure by offering lightweight or offline versions for areas with limited internet connectivity.
@@
Ensure Privacy. Privacy and security were unresolved concerns for many participants. To support adoption under organizational and national constraints, AI systems should: (1) Support on-premise and local AI: allow use without sending code to external servers, enabling companies to keep data within national or internal infrastructure; (2) Enable auditability: ensure that AI-generated code is clearly marked to support inspection and maintain trust.
@@
Simplifying Programming Tasks. Look Beyond Code Writing. To support the needs of all users’ groups, AI should not only assist in code writing but also in code-base navigation, evaluation, and in improving code quality: (1) Codebase navigation: support developers in understanding and exploring the codebase through contextual summaries and links between related components; (2) Project analysis: AI assistants should be able to provide evaluations of the codebase and suggest optimizations for performance, security, or maintainability.
Integrate AI Systems. Participants in all three groups expressed an unmet need for in-IDE AI systems that are tightly connected with internal systems and other engineering tools. Smooth integration with external systems will boost the AI’s utility throughout the SDLC: (1) CI/CD integration: enable the AI to support continuous integration and continuous deployment pipelines, automating build and deployment tasks; (2) Task tracking integration: connect with task management tools, allowing the AI to monitor issue statuses, update progress, and link IDE activities with external sources; (3) Commit management: automate and manage commits, generating commit messages and simplifying the version control process within IDE; (4) Code review support: integrate with code review platforms to assist in reviewing code, and suggesting improvements without leaving the IDE.
Technology Interaction. Streamline In-IDE Interaction. Participants emphasized the importance of minimizing context switching and ensuring smooth interaction between AI and existing IDE features. To support this: (1) Enable embedded interfaces: keep AI interaction within the IDE through integrated chat and in-line code explanations to reduce disruption; (2) Support multimodal input: allow voice commands or device-based control (e.g., wearables) to increase accessibility and flexibility.
Technology Customization. Personalize AI. Developers, especially Churners, seek personalized AI experiences that adapt to their unique needs and workflows. This can be achieved by: (1) Custom setups: allow users to configure the AI assistant, e.g., selecting models, defining interaction methods, or choosing the technology stack; (2) Learning from interactions: let the assistant learn and adjust from user feedback (e.g., declining specific completions).
@@
Simplifying Skill Building. Reduce the Learning Curve. Onboarding and education are key for users to fully utilize AI assistants. All three groups of participants expressed challenges and needs in this regard: (1) User training: offer tutorials and examples to help users interact effectively with the AI; (2) Prompting assistance: implement prompt suggestions and provide users with feedback on their prompts to improve their AI interaction skills.
6. THREATS TO VALIDITY
@@
The prediction market participants were JetBrains employees whose predictions may reflect insider knowledge. While this limits external generalizability, it strengthens ecological validity as these practitioners directly shape product decisions.
@@
External validity. The sample was recruited from a pool of developers using JetBrains products. By focusing on JetBrains’ users but allowing for different AI tools, we aimed to balance consistency with a broad range of experiences, ensuring that our findings are applicable beyond any single tool. However, it may not fully capture the diversity of experiences and perspectives of developers using different IDEs, potentially limiting the findings’ generalizability. Furthermore, our interview sample consists predominantly of senior practitioners, potentially underrepresenting juniors’ needs. Future research should recruit across experience levels and IDE ecosystems to better capture variation in AI tool expectations.
Additionally, the interviews and prediction market were conducted during May-June 2024 and December 2024 - June 2025, respectively, a period characterized by rapid evolution in AI-assisted development tools and changing developer attitudes toward AI integration. The findings reflect developer experiences and expectations during this specific temporal context. Given the fast-paced nature of AI tool development and adoption, developers needs, trust levels, and feature priorities may shift as the technology matures and becomes more ubiquitous. Future research should consider how temporal factors influence the stability of the identified personas and feature preferences over time.
Construct validity. Our thematic analysis followed established qualitative research protocols with iterative coding and consensus-building between two researchers to ensure consistency and reliability. Our goal was to surface previously unexplored differences across adoption stages to inform product strategy, not to establish statistical significance. While our approach cannot statistically validate whether observed thematic patterns represent true group distinctions versus differential reporting rates, this limitation is inherent to exploratory qualitative research with modest sample sizes. The practical value of these insights is demonstrated by their successful translation into actionable features in our prediction market phase, where internal experts and practitioners recognized the feasibility of the features addressing the user needs we identified. Future work could complement our exploratory findings with larger-scale quantitative validation, but the qualitative patterns we uncovered provide essential foundational insights for understanding developer personas in AI tool adoption.
@@
In the second phase, conceptual features derived from interviews were used to formulate prediction market questions. However, these features did not always align well with concrete implementation outcomes. While a structured resolution guide and expert panel were used to mitigate this issue, some interpretive variability may remain, affecting the alignment between predicted outcomes and actual market developments.
7. CONCLUSION
@@
This study examined what different groups of developers want and expect from AI in IDEs, focusing on unmet needs and anticipated features. Through interviews with 35 Adopters, Churners, and Non-Users, we surfaced a structured landscape of user expectations across five areas:Technology Improvements, Programming Tasks, Interaction, Customization, and Skill Building. These highlight the demand for more advanced, reliable, and trustworthy AI experiences within development environments. To assess implementation feasibility, we conducted an internal prediction market with 102 practitioners betting on 9 AI features covering unmet needs. The results show alignment between user needs and practitioner judgment for core implementation and context-aware features. However, proactive behaviors and long-term code maintenance remain both underdelivered and underestimated, pointing to a systemic blind spot. Our findings bridge user expectations and product planning by showing where practitioner foresight aligns, or fails to align, with real-world developer needs. Tool builders should prioritize context-awareness and implementation support today, while investing in proactive and maintenance features for the longer term. Future work should explore how to operationalize unmet users’ needs that are recognized but deprioritized by tool builders.
Acknowledgements.
This work was supported by JetBrains as part of the AI for Software Engineering (AI4SE) collaboration with Delft University of Technology.